Please help! I am really stuck
q^2+6q-7=0

Answer by Alan3354(21583)

q^2+6q-7=0
-------------
That can be factored.
(q+7)*(q-1) = 0
If the product is zero, one of the binomials must be zero.
q+7 = 0
q = -7
------------
q-1 = 0
q = 1